OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that passes through (3, -1) and (-1, 5)? y= - 2y= -3x+ 7 2x+ 3y- 7 = 0 wich answer is it

OpenStudy (hoblos):

it is the 1st answer

OpenStudy (anonymous):

because 5-(-1)/-1-(3)=-3/2 answer 2 th :)

OpenStudy (anonymous):

slope =-3/2

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

because 5-(-1)/-1-(3)=-3/2 answer 2nd :)

OpenStudy (anonymous):

A line passing through two points can be represented in the slope-intercept form as y = mx + b. Where m is the slope and b is the y intercept. But, first we need to find the slope. Slope = difference in y coordinates of the two given points / difference in x coordinates of the two given points. In this case, slope = (-1-5)/(3-(-1)) = -6/4 = -3/2. So, the equation is y = -3x/2 + b. To find b, we should use one of the points on the line (any one of the two really) and plug values for y and x. Let us pick the point (3, -1). We get: -1 = -3(3)/2 + b -1 = -9/2 + b => b = 9/2 - 1 = 7/2 So, the equation is: y = -3x/2 + 7/2

OpenStudy (anonymous):

my answer is true ? :)

OpenStudy (hoblos):

so it is the 1st answer not the 2nd

OpenStudy (anonymous):

hoblos - you getting it right is great. but, helping the others understand to get it is the important thing.